OpenStudy (anonymous):

Which is a solution of the equation? 2x y = 5

OpenStudy (anonymous):

What are you asking? :o

OpenStudy (abb0t):

Is this: \[2x \frac{ dy }{ dx } = 5\]

OpenStudy (anonymous):

2x-y=5

OpenStudy (abb0t):

Are you solving for y?

OpenStudy (anonymous):

If your wanting to know how to make it an equation it would be 2x-y=5.

OpenStudy (abb0t):

1. subtract 2x 2. multiply everything by (-1) to get y and not -y.

OpenStudy (anonymous):

1. Which is a solution of the equation? 2x y = 5 (Points : 4) (7, 3) (2, 1) (1, 2) (3, 1)

OpenStudy (abb0t):

Just plug in the values given.

OpenStudy (abb0t):

when you plug them in, which ever 2 values give you 5, is ur answer.

OpenStudy (anonymous):

2x-y=5 wouldnt it be (1,2)?

OpenStudy (abb0t):

No. (3,1). 2(3)-1 = 5

OpenStudy (abb0t):

Making the solution true.